What is the acceleration due to gravity on the top of Mount Everest? Mount Everest is the highest mountain peak of the world at the height of 8848 m. The value at sea level is 9.80 m s^-2.

Grade:upto college level

1 Answers

Jitender Pal
9 years ago
Sol. Let g’ be the acceleration due to gravity on mount everest. g’ = g (1 – 2h/h) = 9.8 (1 – 17696/6400000) = 9.8 (1 – 0.00276) = 9.773 m/s^2